OpenStudy (anonymous):

Find the indicated function value, if it exists, given f(x) =2-x and g(x)= sqrt3-x. (f+g)(-3)

OpenStudy (anonymous):

Hi! given \(\sf f(x) =2-x\) and \(\sf g(x)= \sqrt{3-x}\) and x=-3 they are asking for \(\sf (f+g)(-3)= f(-3)+g(-3)\) evaluate \(\sf f(-3)+g(-3)\) \(\sf f(-3)+g(-3) =2- (-3) +\sqrt{3-(-3)} \\=2+3+\sqrt{3+3}\\=?\)

OpenStudy (anonymous):

\[5+\sqrt{6} ?\]

OpenStudy (anonymous):

yes that would be right :)